# Analog Quantum Simulation of Coupled Electron-Nuclear Dynamics in Molecules

### v2

@@ IV Conclusion
In this paper, we proposed an analog quantum simulation approach to simulate the quantum dynamics of vibrations and electrons in molecules. Our approach uses a pre-BO wavefunction ansatz, which converges to the exact solution of the non-relativistic time-dependent Schrödinger equation with increasing electronic basis set size and nuclear Taylor expansion order. Furthermore, our approach can be extended to describe intersystem crossing and ionization without additional resources, and can be applied to quantum simulations of electron-phonon coupling models for solids. In contrast, most previously proposed approaches for the simulation of molecular dynamics employ the GBOA, which artificially truncates the electronic state basis. Others use a real-space, first quantization approach, which includes molecular rotations that complicate the interpretability of the wavefunction.
For equivalent descriptions of vibronic dynamics, our approach shows clear scaling advantages over all proposals to date when the quantum resources, number of operations, and number of pre-calculations are considered. This work makes use of existing tools in analog quantum simulation, and model systems such as the example presented in this work can be readily implemented on existing quantum hardware such as trapped ion quantum computers with realistic noise effects. With improvements in quantum hardware and quantum control, we expect that our approach will demonstrate an early quantum advantage for the simulation of quantum chemistry.

### v3

@@ IV Conclusion
In this paper, we proposed a quantum simulation approach with an analog mapping of nuclear degrees of freedom to simulate the quantum dynamics of vibrations and electrons in molecules. We presented an emulation of the digital-analog implementation of our approach on a trapped-ion device, using a simple vibronic dynamics model system resembling realistic chemical processes. Our approach uses a pre-BO wavefunction ansatz, which converges to the exact solution of the non-relativistic time-dependent Schrödinger equation with increasing electronic basis set size and nuclear Taylor expansion order. Furthermore, our approach can be extended to describe intersystem crossing and ionization without additional resources, and can be applied to quantum simulations of electron-phonon coupling models for solids. In contrast, most previously proposed approaches for the simulation of molecular dynamics employ the BO framework, which often involves truncation in the electronic basis in dynamics simulation and/or electronic structure calculations for the BO states and couplings. Others use a real-space, first quantization approach, which includes molecular rotations that complicate the interpretability of the wavefunction.
For equivalent descriptions of vibronic dynamics, our approach shows clear scaling advantages over existing quantum algorithms when the accuracy, quantum resources, number of operations, and number of pre-calculations are considered. However, true advantages of our approach for realistic molecules require careful consideration of potential sources of error such as Trotterization, noise, and Taylor expansion truncation, as well as comparison to classical computing approximations. Model systems such as the example presented in this work can be readily implemented on existing quantum hardware such as trapped ion quantum computers with realistic experimental noise. With improvements in quantum hardware and quantum control, we expect that our approach will demonstrate an early quantum advantage for the simulation of quantum chemistry.